Vinegar is a solution of acetic acid in water. If a 325
mL
bottle of distilled vinegar contains 28.6
mL
of acetic acid, what is the volume percent (v/v) of the solution?
Express your answer to three significant figures.
1
Expert's answer
2019-11-25T07:25:45-0500

volume percent (v/v) = "\frac{28.6}{325} \times 100\% = 88.0\%"